Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
Nenci - 6:57 Sunday 01 December 2019 (47824) Print this report
Operator Report - Night shift

ITF stable locked in Science Mode for all the shift, BNS Range ~49 Mpc.

ISYS -  BsX_PZT is still increasing (picture 1 attached). "Oncall events" for details.

Air Conditioning / INJ_Area - In the attached plot 2 you see a trend of HVAC..INJ_LLR_TE, the temperature seems to be stabilized at 23 C instead 22.5 C.

Guard Tour (UTC)
23:45 - 00:20
01:40 - 02:09
03:30 - 04:01
05:02 - 05:34

ISYS
(01-12-2019 00:00 - ) Upon channel BsX_PZT on DMS will reach the critical value of 7,5 / 8 ISYS OnCall must contacted.

ISYS
(30-11-2019 23:00 - 30-11-2019 23:15) From remote
Status: Ended
Description: BsX_PZT near to the critical value, therefore i contacted the expert in order to check the situation. The signal was following the temperature in the Injection area that was not still stabilized, will be monitored during the night. The intervention onsite will be needed if the value will increase again and go up ~ 7,5 / 8.

Images attached to this report
Comments to this report:
chiummo, Gherardini - 10:51 Sunday 01 December 2019 (47826) Print this report

Due to the HVAC failure, a large temperature excursion occurred in the laser lab. This, among other effects, has triggered a drift of the beam pointing compensated by the BPC loop with large corrections, especially on the horinzontal degrees of freedom (see first plot).

Unfortunately, after the fixing of the HVAC, the temperatures took a long time to stabilize, and are still experiencing a long transient. Nevertheless, the drift of the beam pointing had become too large to be dealt with by the correction dynamics, so I come on site to decrease the needed corrections of the BPC by acting on the usual M6-M8 picomotors. The operation has been successful (see second plot), but fine tuning will be needed once that the steady state of the temperature will be achieved at the right values.

The interferometer was able to relock up to LN3 and was set in science mode after the operation.

Images attached to this comment
Search Help
×

Warning

×